Big plane. A six-passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind, then what is the wind speed?
Answer by richwmiller(17219)

(180+w)*t=7
(180-w)*t=5
t=7/(180+w)
t=5/(180-w)
7/(180+w)=5/(180-w)
w=30